OpenStudy (anonymous):

What is the lateral area of a pyramid with base edges 5ft and surface area 55ft^2?

OpenStudy (dumbcow):

lateral area = 1/2 *perimeter of base * slant height total surface area = lateral area + AreaBase thus lateral area = 55 - 5^2 = 55-25 = 30

OpenStudy (anonymous):

Surface area = Lateral area + Base area therefore 55ft2 = Lateral Area + 5ft x 5ft Lateral Area = 30ft2